Determine if the following lines are parallel, skew or intersecting.




⟨x,y,z⟩=⟨2,1,2⟩+⟨3,5,2⟩t and



⟨x,y,z⟩=⟨2,5,2 ⟩+⟨3,3,2⟩s



*If parallel (type the number 1).



*If intersecting (type the number 2).



*If skew (type the number 3).


1
Expert's answer
2022-03-14T20:07:38-0400

"<x,y,z>=<2,1,2>+<3,5,2>t\\
<x,y,z>=<2,5,2>+<3,3,2>s\\
1. \vec{a_1}=(3,5,2)\\
 \vec{a_2}=(3,3,2)\\
\frac{3}{3}\neq\frac{5}{3}\neq\frac{2}{2}"

lines are not parallel

2.

"A(2,1,2), B(2,5,2)\\ \overrightarrow{AB}=(2-2,5-1,2-2)=(0,4,0)\\
\overrightarrow{AB}\cdot\vec{a_1}\cdot\vec{a_2}=\\
=\begin{vmatrix}
 0 & 4&0 \\
 3 & 5&2\\
3&3&2
\end{vmatrix}=0\cdot5\cdot2+4\cdot2\cdot3+0\cdot3\cdot3-\\
-0\cdot5\cdot3-3\cdot4\cdot2-3\cdot2\cdot0=24-24=0"

lines are intersecting

*If intersecting (type the number 2).